# Premerger phenomena in neutron-star binary coalescences

### v1

@@ I Introduction
Massive stars that have exhausted their fuel reservoir eventually collapse, as the outward pressures no longer resist the inward gravitational pull. The brilliant supernovae that accompany these collapses rip away most of the outer layers of material, though the inner layers become more and more compressed. Subatomic fermions provide a last bastion of resistance against this compression through their degeneracy pressures because of the Pauli exclusion principle. The strongest such pressure is from neutrons, and thus neutron stars represent some of the most compact objects in the universe, behind only black holes and the hypothetical quark stars. Neutron-star core densities likely reach supranuclear levels [$\sim 10^{15}\text{ g/cm}^{3}$; 286, 283, 235] because their radii are limited to $\gtrsim 10$ km. Studying their properties thus allows, for example, for tests of the “low-energy” limit of quantum chromodynamics with baryonic degrees of freedom and general relativity (GR). The exact chemical makeup of a star defines its equation of state (EOS), determining the particulars of which constitutes one of the open problems in high-energy astrophysics.
@@
Neutron-star mergers are some of the brightest events in the universe. A crescendo of gravitational waves (GWs) are released during the final moments of inspiral culminating in a “chirp” and subsequent ringdown of the hyperactive remnant left behind at the crash site, being either a black hole or another, heavier neutron star [e.g. 350, 349]. In the latter case the object will likely be short-lived: it may maintain a state of metastability despite having a mass larger than the maximum (i.e. Tolman-Oppenheimer-Volkoff limit) applying to an “ordinary” neutron star if supported by centrifugal [135], thermal [200], or magnetic [374] pressures.
@@
Much of the fireworks come shortly after the actual coalescence in the form of gamma-ray bursts (GRBs), primarily of the short variety with prompt durations less than two seconds though there is evidence that some mergers can also fuel long bursts [e.g. 399, 238]. Through either a fireball-like (neutrino-anti-neutrino) or Poynting-like process a relativistic jet is collimated by the remnant, providing the impetus for a wealth of emissions that are both luminous (isotropic energies often reach $\sim 10^{53}$ erg) and broadband (spanning $\sim 20$ magnitudes in frequency space via long-lived afterglows). The multimessenger event GW170817/GRB170817A [7, 163] has not only firmly established the merger-GRB connection but has been used to place strong constraints on the nature of matter in extreme environments and hypothetical departures from GR in the ultraviolet [e.g. 8, 84, 120].
@@
A range of luminous and sometimes multiband events can also take place in the moments leading up to a binary coalescence. One of the main subjects of this Review pertains to such precursors: a fraction [$\lesssim 5\%$; 417] of merger-driven GRBs show statistically-significant gamma-ray flashes before the main GRB. While other types of precursors may be released premerger (covered in Sec. VII), we typically write precursor to specifically refer to these first-round gamma flares. Indeed, many reviews are devoted to the topic of merger phenomena, involving numerical simulations and observations of inspiral, remnant dynamics, and the subsequent forming of relativistic jets [see 134, 43, 347, 73, 320, 90, 339, 206, for instance]. Comparatively little attention has been given towards a complete description of what can be (and has been!) learned just prior to merger where such precursors may be expected. Our goal is to fill this gap, motivated in part by recent observational campaigns devoted to searching for precursor flares [e.g. 96, 420, 111] and the ushering in of the LIGO-Virgo-KAGRA (LVK) collaboration’s O4 science run. In the spirit of the burgeoning field of multimessenger astrophysics, we also cover the theory of tides and premerger GWs in detail in the hope that a reader not interested in gamma-ray precursors may still find value.
@@
Precursor flashes give promise for a plethora of information about fundamental physics, naturally complementing that associated with merger and postmerger phenomena. A premerger precursor could be used, for instance, to improve sky localization for an impending collision if nothing else [see, e.g. 94]. One key difference though is that premerger objects are likely cold, and thus arguably give a better handle on the EOS as the impact of $\gg$ MeV temperatures and a littered environment (e.g. from dynamical ejecta) do not require disentangling. Precursors are observationally diverse. They have waiting times — relative to the main GRB (not merger!) — spanning $\sim 30$ ms to $\sim$20 s. Their durations and luminosities span a commensurate number of magnitudes, with spectra varying from being almost a perfect blackbody to being highly non-thermal. Such an assortment of characteristics makes a strong case that there may be subpopulations of such precursors fuelled by different means [e.g. 417]. By dividing our discussion on precursors into observational (Sec. V) and theoretical (Sec. VI) elements, we hope to build a full picture of these rich systems.
@@
Premerger precursors are primarily thought to be caused by one of two means. The first involves direct, electromagnetic interactions between the two stars [e.g. 175, 287, 419, 415]. For instance, if the inspiralling constituents have dipole moments which are anti-aligned with respect to each other, ample reconnection can take place as the field lines get more and more entwined as the orbit decays. To get large luminosities one would expect the flares to be emitted very close to merger as the dipole fields die off like the cube of distance. One may anticipate relatively short waiting times (see Sec. V) in this case therefore, unless magnetars are involved in the merger. Spectral and other considerations of certain precursors however give reason to suspect that some mergers do indeed contain a magnetar [e.g. 118, 381, 428], despite the fact that the GW inspiral time exceeds the characteristic Ohmic-decay timescale by orders of magnitude (see Sec. II.3).
@@
The second ignition mechanism involves tidal fields. Tides can not only deform the star geometrically (“equilibrium tide”) but also excite internal fluid motions (“dynamical tides”). These internal motions are characterised through a set of quasi-normal modes (QNMs). These QNMs, which come in a few families (e.g. $g$-modes) that we properly introduce in Secs. IV.1.3 and VI.4 though sprinkle information about throughout, can be driven to large amplitudes when coming into resonance with the dynamical tidal field. This tidal driving can manifest gravitationally, through a dephasing of the gravitational waveform, and electromagnetically, by liberating potential energy a star whose crustal layers succumb to the high-amplitude pulsations. Ruptures and subsequent energy release models are akin to some of those considered for magnetar flares [see, e.g. 393, 394], though in the premerger context the cause of crustal failure is QNM resonances rather than the gradual build-up of mechanical stress from the secular evolution of a superstrong field.
@@
Galactic neutron stars exhibit a variety of multiwavelength phenomena, the varying characteristics of which has invited rather detailed classification schemes over the years, with stars being grouped into categories such as recycled, radio, or X-ray pulsars [e.g. 131]. We describe how precursors, and inspiral lead-up more generally, can inform us about each of these pathways. This Review is thus roughly divided into two halves: that devoted to GWs and that to electromagnetic precursors, described in more detail below.
@@ II.1 Equations of state
Neutron stars exhibit an astonishing range of multiwavelength phenomena, from steady radio pulsations to rare storms of powerful soft-gamma flares [e.g. 285]. From a nuclear physics perspective however, it is thought that neutron-star matter comes into beta equilibrium shortly after birth [191]. This implies that their cores are described by a unique EOS moderated by small, thermodynamic perturbations. An important exception regards stars born out of merger events [19], which can reach $\gg$ MeV temperatures where new particle production pathways are opened. For mature stars a barotropic EOS is typically used though, where the hydrostatic pressure $p$ is just a function of the rest-mass density, $\rho$. This latter quantity is distinct from the energy density $\epsilon$, though sometimes in the literature one can find “EOS” of the form $p=p(\epsilon)$; see Section 2 of O’Boyle et al. 282 for a discussion.
Many different EOS families have been considered in the literature [286, 235], primarily owing to our collective ignorance regarding the behaviour of matter at supra- and hypernuclear densities at low temperatures. The EOS families we consider in an effort to make this Review self-contained — mass–radius curves of which are displayed in Figure 2 — cover stars with purely $nep\mu$ compositions to those exhibiting phase transitions to free-quark regimes [i.e. the hybrid stars, like ALF2; 18]. Although some of the considered EOS cannot support the heaviest (confirmed [^1] For an up-to-date catalogue of neutron-star mass measurements, at least where there are no substantial systematics due to mass transfer or mass loss, see https://www3.mpifr-bonn.mpg.de/staff/pfreire/neutronstar_masses.html. There are supposedly heavier “spiders”, though these are highly dynamical and also rotating rapidly.) neutron star observed to-date, namely PSR J0740+6620 with a mass of $\sim 2.08^{+0.07}_{-0.07}M_{\odot}$, we consider them for completeness (especially those that pass right through the heart of GW 170817 contours). Another object worth mentioning in this landscape is HESS J1731-347, with a mass of $\gtrsim 0.8M_{\odot}$ and $R\approx 10.4$ km [125]; if indeed this was representative of the lower limit to the possible neutron star mass, it would impose serious restrictions to the EOS and formation mechanisms. J1731 measurements would, funnily enough, be totally consistent with the dashed-curve EOS in Fig. 2 that do not reach $M\gtrsim 2M_{\odot}$; see Figure 5 in Ofengeim et al. 284.
@@
Thermal and inviscid effects in a mature neutron star, though crucial for modelling buoyancy-restored oscillations (i.e. $g$-modes; see Sec. IV.1.3) especially in the crust (see Sec. III.1), are often not present in many of the “standard” EOS tables, as the in situ integrals are notoriously difficult to calculate except in some perturbative sense (e.g. including a $\propto T^{2}$ thermal pressure contribution). The “general purpose EOS” (following the naming of the catalogue) available on the CompOSE catalogue [^2] These and many other EOS can be obtained in a tabulated form via the comprehensive CompOSE [403] repository: https://compose.obspm.fr/home/. provide such information. Although a few general purpose EOS do extend to a lower cutoff at $10^{7}$ K (e.g. SLy4 and APR), many of them have their tables truncated at minima of $10^{9}$ K (still much higher than expected of a mature star seconds before merger; though cf. Sec. IV.5) since lower temperatures are well-modelled by cold EOS with some perturbative corrections, as described above.
The impact of realistic thermal profiles on the $g$-spectra in this perturbative sense has been detailed by Krüger et al. 218 and by Kuan et al. 226 with an independent code, finding consistent results. As found by Kuan et al. 222, the $g$-spectra are linked to the “slope” of the EOS $M$ vs. $R$ curve, and are especially sensitive to central densities in the range of 1–3 $\rho_{0}$ for nuclear saturation value $\rho_{0}\approx 2\times 10^{14}\text{ g cm}^{-3}$ [236]. The slope of the $M$$R$ curve impacts on the universality of some relations between bulk properties and dominant peaks in postmerger waveforms [322]. Kuan et al. 222 also found that $g$-mode scaling relations can be grouped according to EOS, with hadronic ones that can support heavy stars (solid lines in Fig. 2), those that cannot (dashed), and EOS involving phase transitions leading to either hyperon condensation or quark deconfinement (dash-dotted). These three families each abide by a different set of “asteroseismological relations” that can, in principle, be distinguished by observations.
Detailed analyses of the EOS population, their historical development, together with astrophysical applications and constraints are found in Lattimer and Prakash 236, Oertel et al. 283, Özel and Freire 286, Lattimer 235. Observationally, masses can be constrained from a variety of methods, ranging from orbital timing of binaries to the phased-resolved tracking of thermal hotspots atop an isolated star to GWs. Observations of the moments of inertia can also be made from periastron advance in binaries [168]. We now turn to describing a few important EOS families.
@@ II.1.1 WFF family
The WFF family (WFF1-3) was developed in the late 1980s [425] based on realistic nucleon-nucleon interactions calculated using many-body methods. This family of EOS treats nucleons as non-relativistic particles, but does utilise the Argonne v14 (so-named because it uses 14 operator components to model nucleon-nucleon interactions) and Urbana vII potentials which account for three-nucleon forces. These two potentials are essential for the accurate modelling the physics of nuclear forces at high densities. The family is generally considered to be a softer EOS, especially WFF1. According to the WFF family of EOS, the maximum mass of a neutron star typically ranges up to ${}2.2M_{\odot}$ (WFF1-2) in agreement with the aforementioned observations. Still, recent results from GW170817 and the Neutron Star Interior Composition Explorer (NICER) suggest a stiffer EOS, thus WFF1 is excluded as being too soft; it is for this reason we have excluded it from Figure 2. Despite these limitations, many works still use the WFF family of EOS as representatives of softer EOS. The WFF family has improved over time by including additional interactions and so on, with many modern EOS stemming from it [see, e.g. 426].
@@ II.1.2 APR family
The APR EOS is named for the authors Akmal, Pandharipande, and Ravenhall who developed it in the late 1990s [15]. It is widely adopted, despite the fact that at high densities relativistic corrections are large and using the non-relativistic Schrödinger equation at $\rho\sim\text{several}\times\rho_{0}$ is questionable, as it meets the maximum mass observational limit in the static limit (and beyond to $\sim 2.3M_{\odot}$), models the cooling evolution (implicitly via proton-fraction predictions and beta-decay rates in superfluids) as observed in young and old neutron stars [see, e.g. 37], and can meet the constraints set by GW observations of neutron-star mergers; it is not necessarily unique in these respects though. The APR EOS is derived from a microscopic nuclear many-body theory using the variational chain summation method. It predicts a phase transition from pure nucleonic matter to a state including nucleons and a neutral pion condensation at high densities [342] and incorporates realistic nucleon-nucleon interactions based on the Argonne v18 potential (which, as per the discussion above, includes 18 interaction operators). In addition to two-body interactions, it includes three-body forces using the Urbana IX potential: these are essential for providing repulsive interactions at high densities. It is through these repulsive interactions that one may construct heavier neutron stars. For up-to-date discussions, see Schneider et al. 342, Burgio et al. 72.

### v2

@@ I Introduction
Massive stars that have exhausted their fuel reservoir eventually collapse, as the outward pressures no longer resist the inward gravitational pull. The brilliant supernovae that accompany these collapses rip away most of the outer layers of material, though the inner layers become more and more compressed. Subatomic fermions provide a last bastion of resistance against this compression through their degeneracy pressures because of the Pauli exclusion principle. The strongest such pressure is from neutrons, and thus neutron stars represent some of the most compact objects in the universe, behind only black holes and the hypothetical quark stars. Neutron-star core densities likely reach supranuclear levels [$\sim 10^{15}\text{ g/cm}^{3}$; 292, 289, 241] because their stellar radii are of order $\sim 10$ km. Studying their properties thus enables, for instance, tests of the “low-energy” limit of quantum chromodynamics with baryonic degrees of freedom and general relativity (GR). The exact chemical makeup of a star defines its equation of state (EOS), determining the particulars of which constitutes one of the open problems in high-energy astrophysics.
@@
Neutron-star mergers are some of the brightest events in the universe. A crescendo of gravitational waves (GWs) are released during the final moments of inspiral culminating in a “chirp” and subsequent ringdown of the hyperactive remnant left behind at the crash site, being either a black hole or another, heavier neutron star [e.g. 357, 356]. In the latter case the object will likely be short-lived: it may maintain a state of metastability despite having a mass larger than the maximum (i.e. Tolman-Oppenheimer-Volkoff limit) applying to an “ordinary” neutron star if supported by centrifugal [136], thermal [205], or magnetic [379] pressures.
@@
Much of the fireworks come shortly after the actual coalescence in the form of gamma-ray bursts (GRBs), primarily of the short variety with prompt durations less than two seconds though there is evidence that some mergers can also fuel long bursts [e.g. 404, 244]. Through either a fireball-like (neutrino-anti-neutrino) or Poynting-like process a relativistic jet is collimated by the remnant, providing the impetus for a wealth of emissions that are both luminous (isotropic energies often reach $\sim 10^{53}$ erg) and broadband (spanning $\sim 20$ magnitudes in frequency space via long-lived afterglows). The multimessenger event GW170817/GRB170817A [7, 165] has not only firmly established the merger-GRB connection but has been used to place strong constraints on the nature of matter in extreme environments and hypothetical departures from GR in the ultraviolet [e.g. 8, 85, 121].
@@
A range of luminous and sometimes multiband events can also take place in the moments leading up to a binary coalescence. One of the main subjects of this Review pertains to such precursors: a fraction [$\lesssim 5\%$; 422] of merger-driven GRBs show statistically-significant gamma-ray flashes before the main GRB. While other types of precursors may be released premerger (covered in Sec. VII), we typically write precursor to specifically refer to these first-round gamma flares. Indeed, many reviews are devoted to the topic of merger phenomena, involving numerical simulations and observations of inspiral, remnant dynamics, and the subsequent forming of relativistic jets [see 135, 44, 354, 74, 325, 91, 345, 211, for instance]. Comparatively little attention has been given towards a complete description of what can be (and has been!) learned just prior to merger where such precursors may be expected. Our goal is to fill this gap, motivated in part by recent observational campaigns devoted to searching for precursor flares [e.g. 97, 425, 112] and the ushering in of the LIGO-Virgo-KAGRA (LVK) collaboration’s O4 science run. In the spirit of the burgeoning field of multimessenger astrophysics, we also cover the theory of tides and premerger GWs in detail in the hope that a reader not interested in gamma-ray precursors may still find value.
@@
Precursor flashes give promise for a plethora of information about fundamental physics, naturally complementing that associated with merger and postmerger phenomena. A premerger precursor could be used, for instance, to improve sky localization for an impending collision if nothing else [see, e.g. 95]. One key difference though is that premerger objects are likely cold, and thus arguably give a better handle on the EOS as the impact of $\gg$ MeV temperatures and a littered environment (e.g. from dynamical ejecta) do not require disentangling. Precursors are observationally diverse. They have waiting times — relative to the main GRB (not merger!) — spanning $\sim 30$ ms to $\sim$20 s. Their durations and luminosities span a commensurate number of magnitudes, with spectra varying from being almost a perfect blackbody to being highly non-thermal. Such an assortment of characteristics makes a strong case that there may be subpopulations of such precursors fuelled by different means [e.g. 422]. By dividing our discussion on precursors into observational (Sec. V) and theoretical (Sec. VI) elements, we hope to build a full picture of these rich systems.
@@
Premerger precursors are primarily thought to be caused by one of two means. The first involves direct, electromagnetic interactions between the two stars [e.g. 177, 293, 424, 420]. For instance, if the inspiralling constituents have dipole moments which are anti-aligned with respect to each other, ample reconnection can take place as the field lines get more and more entwined as the orbit decays. To get large luminosities one would expect the flares to be emitted very close to merger as the dipole fields die off like the cube of distance. One may anticipate relatively short waiting times (see Sec. V) in this case therefore, unless magnetars are involved in the merger. Spectral and other considerations of certain precursors however give reason to suspect that some mergers do indeed contain a magnetar [e.g. 119, 386, 433], despite the fact that the GW inspiral time exceeds the characteristic Ohmic-decay timescale by orders of magnitude (see Sec. II.3).
@@
The second ignition mechanism involves tidal fields. Tides can not only deform the star geometrically (“equilibrium tide”) but also excite internal fluid motions (“dynamical tides”). These internal motions are characterised through a set of quasi-normal modes (QNMs). These QNMs, which come in a few families (e.g. $g$-modes) that we properly introduce in Secs. IV.1.3 and VI.4 though sprinkle information about throughout, can be driven to large amplitudes when coming into resonance with the dynamical tidal field. This tidal driving can manifest gravitationally, through a dephasing of the gravitational waveform, and electromagnetically, by liberating potential energy a star whose crustal layers succumb to the high-amplitude pulsations. Ruptures and subsequent energy release models are akin to some of those considered for magnetar flares [see, e.g. 398, 399], though in the premerger context the cause of crustal failure is QNM resonances rather than the gradual build-up of mechanical stress from the secular evolution of a superstrong field.
@@
Galactic neutron stars exhibit a variety of multiwavelength phenomena, the varying characteristics of which has invited rather detailed classification schemes over the years, with stars being grouped into categories such as recycled, radio, or X-ray pulsars [e.g. 132]. We describe how precursors, and inspiral lead-up more generally, can inform us about each of these pathways. This Review is thus roughly divided into two halves: that devoted to GWs and that to electromagnetic precursors, described in more detail below.
@@ II.1 Equations of state
Neutron stars exhibit an astonishing range of multiwavelength phenomena, from steady radio pulsing to rare storms of powerful soft-gamma flares [e.g. 291]. From a nuclear physics perspective however, it is thought that neutron-star matter comes into beta equilibrium shortly after birth [194]. This implies that their cores are described by a unique EOS moderated by small, thermodynamic perturbations. Proto-stars, especially those born in merger events [20], may be exempt however as new particle production pathways can be opened at super-MeV temperatures. For mature stars a barotropic EOS is typically used though, where the hydrostatic pressure $p$ is just a function of the rest-mass density, $\rho$. This latter quantity is distinct from the energy density $\epsilon$, though sometimes in the literature one can find “EOS” of the form $p=p(\epsilon)$; see Section 2 of O’Boyle et al. 288 for a discussion.
Many EOS families have been considered in the literature [292, 241], primarily owing to our collective ignorance regarding the behaviour of matter at supra- and hypernuclear densities at low temperatures. The EOS families we consider in an effort to make this Review self-contained — mass–radius curves of which are displayed in Figure 2 — cover stars with purely $nep\mu$ compositions to those exhibiting phase transitions to free-quark regimes [i.e. the hybrid stars, like ALF2; 18]. Although some of the considered EOS cannot support the heaviest (confirmed [^1] For an up-to-date catalogue of neutron-star mass measurements, at least where there are no substantial systematics due to mass transfer or mass loss, see https://www3.mpifr-bonn.mpg.de/staff/pfreire/neutronstar_masses.html. There are supposedly heavier “spiders”, though these are highly dynamical and also rotating rapidly.) neutron star observed to-date, namely PSR J0740+6620 with a mass of $\sim 2.08^{+0.07}_{-0.07}M_{\odot}$, we consider them for completeness (especially those that pass right through the heart of GW 170817 contours). Another object worth mentioning in this landscape is HESS J1731-347, with a mass of $\gtrsim 0.8M_{\odot}$ and $R\approx 10.4$ km [126]; if indeed this was representative of the lower limit to the possible neutron star mass, it would impose serious restrictions to the EOS and formation mechanisms. J1731 measurements would, funnily enough, be totally consistent with the dashed-curve EOS in Fig. 2 that do not reach $M\gtrsim 2M_{\odot}$; see Figure 5 in Ofengeim et al. 290.
@@
Thermal and inviscid effects in a mature neutron star, though crucial for modelling buoyancy-restored oscillations (i.e. $g$-modes; see Sec. IV.1.3) especially in the crust (see Sec. III.1), are often not present in many of the “standard” EOS tables because the effects are thought to be safely negligible below $\sim 0.1$ MeV where the chemical potential is tiny. Such corrections can be handled with a $\propto T^{2}$ thermal pressure contribution. The “general purpose EOS” (following the naming of the catalogue) available on the CompOSE catalogue [^2] These and many other EOS can be obtained in a tabulated form at [408]: https://compose.obspm.fr/home/. provide such information. Although a few general purpose EOS do extend to a lower cutoff at $10^{7}$ K (e.g. SLy4 and APR), many of them have their tables truncated at minima of $10^{9}$ K (still much higher than expected of a mature star seconds before merger; though cf. Sec. IV.5) since lower temperatures are well-modelled by cold EOS with some perturbative corrections, as described above.
The impact of realistic thermal profiles on the $g$-spectra in this perturbative sense has been detailed by Krüger et al. 224 and by Kuan et al. 232 with an independent code, finding consistent results. As found by Kuan et al. 228, the $g$-spectra are linked to the “slope” of the EOS $M$ vs. $R$ curve, and are especially sensitive to central densities in the range of 1–3 $\rho_{0}$ for nuclear saturation value $\rho_{0}\approx 2\times 10^{14}\text{ g cm}^{-3}$ [242]. The same slope impacts on the universality of some relations between bulk properties and dominant peaks in postmerger waveforms [327]. Kuan et al. 228 also found that $g$-mode scaling relations can be grouped according to EOS, with hadronic ones that can support heavy stars (solid lines in Fig. 2), those that cannot (dashed), and EOS involving phase transitions leading to either hyperon nucleation or quark deconfinement (dash-dotted); we refer the reader to Sedrakian et al. 350 for a review on hyperonisation. These three families each abide by a different set of “asteroseismological relations” that can, in principle, be distinguished by observations.
Detailed analyses of the EOS population, their historical development, together with astrophysical applications and constraints are found in Lattimer and Prakash 242, Oertel et al. 289, Özel and Freire 292, Lattimer 241. Observationally, masses can be constrained from a variety of methods, ranging from orbital timing of binaries to the phased-resolved tracking of thermal hotspots atop an isolated star to GWs. Observations of the moments of inertia can also be made from periastron advance in binaries [170]. We now turn to describing a few EOS families. Note that some of the EOS appearing in Fig. 2 are not covered below despite their consistency with data, though could adequately be described as “relativistic models.” For example, MPA and ENG are based on Dirac-Brueckner-Hartree-Fock models for purely nucleonic matter, while GNH(3) is a relativistic mean-field EOS that also contains hyperonic matter. In one of the considered EOS (ALF2) a hadron-quark hybrid core is present, described through the MIT bag model.
@@ II.1.1 WFF family
The WFF family (WFF1-3) was developed in the late 1980s [430] based on realistic nucleon-nucleon interactions calculated using many-body methods. This family treats nucleons as non-relativistic particles, utilising the Argonne v14 (so-named because it uses 14 operator components to model nucleon-nucleon interactions) and Urbana vII potentials which account for three-nucleon forces. The family is generally considered to be a softer EOS, especially WFF1. According to the WFF family of EOS, the maximum mass of a neutron star typically ranges up to ${}2.2M_{\odot}$ (WFF1-2) in agreement with the aforementioned observations. Still, recent results from GW170817 and the Neutron Star Interior Composition Explorer (NICER) suggest a stiffer EOS, thus WFF1 is excluded as being too soft; it is for this reason we have excluded it from Figure 2. Despite these limitations, many works still use the WFF family of EOS as representatives of softer EOS. The WFF family has improved over time by including additional interactions and so on, with many modern EOS stemming from it [see, e.g. 431].
@@ II.1.2 APR family
The APR EOS is named for the authors Akmal, Pandharipande, and Ravenhall who developed it in the late 1990s [15]. It is widely adopted, despite the fact that at high densities relativistic corrections are large and using the non-relativistic Schrödinger equation at $\rho\sim\text{several}\times\rho_{0}$ is questionable, as it meets the maximum mass observational limit in the static limit (and beyond to $\sim 2.3M_{\odot}$), models the cooling evolution (implicitly via proton-fraction predictions and beta-decay rates in superfluids) as observed in young and old neutron stars [see, e.g. 38], and can meet the constraints set by GW observations of neutron-star mergers; it is not necessarily unique in these respects though. The APR EOS is derived from a microscopic nuclear many-body theory using the variational chain summation method. It predicts a phase transition from pure nucleonic matter to a state including nucleons and a neutral pion condensation at high densities [348] and incorporates realistic nucleon-nucleon interactions based on the Argonne v18 potential (which, as per the discussion above, includes 18 interaction operators). In addition to two-body interactions, it includes three-body forces using the Urbana IX potential: these are essential for providing repulsive interactions at high densities. It is through these repulsive interactions that one may construct heavier neutron stars. For up-to-date discussions, see Schneider et al. 348, Burgio et al. 73.